Hi. I'm looking for the quadratic equation that describes the following three sets of coordinates: (5,19) (75,75) (150, 101). Please help. I have no idea where to begin. Thanks!!!
Answer by lwsshak3(11628)

Hi. I'm looking for the quadratic equation that describes the following three sets of coordinates: (5,19) (75,75) (150, 101). Please help. I have no idea where to begin.
**
I will show you how to do it, but you must work it out yourself:
Remember the quadratic equation in the form: ax^2+bx+c=0
It can also be written as: ax^2+bx+c=y
y is set to zero when finding the roots or zeros.
..
With the given 3 sets of (x,y) coordinates, you need to find a, b, and c, with the following 3 equations:
(5^2)a+5b+c=19
(75^2)a+75b+c=75
(150^2)a+150b+c=101
after finding a, b, and c, plug them into the basic quadratic form of equation: y=ax^2+bx+c
..
The numbers are such that finding a, b and c, would require a whole lot of careful algebra work beyond what I would be willing to do, but this is how you do it if you are willing to try.